# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v2

@@ 3.3 Line blending to align similar linestrings with local reference
So far we have only considered line blending where the reference linestring has a single flow. Due to the noisiness of the empirical trajectories, this is insufficient to reach a minimal flow map. So we allow the reference to be a sequence of $k$ connected edges, each with their own flow. If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings. Due to computational limitations, we retain that candidates be single edges with single flow values. The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
@@
We have described the situation for blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges). The next step to determine the priority for line blending for a set of reference linestrings. We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only. These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them. The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except for where $f_{1}=f_{2}$. In the situation of equal flow values, then we designate the longer linestring to be the reference.
@@
The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$. The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$ and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings. In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e. all paths composed of connected, unique edges) with $k$-edges. In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path. The function ${\rm Edges}(\cdot)$ extracts the edges from a $k$-edge path. In Step 5, we sort the flow linestrings, in descending order of their flow values and length. In Steps 6–8, we construct a maximal set of non-overlapping $k$-edge paths $\mathcal{K}^{*}$. We initialise $\mathcal{K}^{*}$ to contain the first path. We iterate through $\mathcal{K}$ and, if this path in $\mathcal{K}$ does not overlap the current $\mathcal{K}^{*}$, then we add it to $\mathcal{K}^{*}$. For $k=1$, we bypass Steps 2–8. In Steps 9–18, we step through the ${\bm{k}}_{i}^{*}\in\mathcal{K}^{*}$, starting from the linestring with the highest flow and length. In Steps 11–13, if the current flow linestring ${\bm{k}}_{i}^{*}$ has no incident edges in the candidate linestrings $\mathcal{F}_{\rm cand}$, then we set it to be a reference linestring. In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1). In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$ and $\mathcal{F}_{\rm ct}$. In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets, since the reference linestring set is already computed as $\mathcal{F}_{\rm ref}$.
@@
6:     $f_{i}$ := weighted mean of $k$ flows, weight := len(Edges(${\bm{k}}_{i}$))
@@
13:     if (${\rm Edges}({\bm{k}}_{i}^{*})\notin\mathcal{F}_{\rm cand}$then
@@
15:         Search candidate linestrings $\mathcal{F}_{{\rm cand}}^{*}:=\{{\bm{f}}\in\mathcal{F}\backslash({\rm Edges}(\mathcal{F}_{\rm ref})\cup\{{\bm{k}}_{\rm ref}\}\cup\mathcal{F}_{\rm cand}):{\bm{f}}\subseteq$
16:    ${\tt ST\_BUFFER}({\bm{k}}_{\rm ref},\varepsilon),f\leq f_{\rm ref}\}$
17:         if $\mathcal{F}_{\rm cand}^{*}\neq\{\}$ then
18:              Update $\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm ref}\cup\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm cand}^{*}$
19:              Search touches $\mathcal{F}_{{\rm ct}}:=\{{\bm{f}}\in\mathcal{F}\backslash\{{\rm Edges}(\mathcal{F}_{\rm ref})\ \cup\mathcal{F}_{\rm cand}\}:{\tt ST\_TOUCHES}(\mathcal{F}_{{\rm cand}}^{*},{\bm{f}})\}$
20:              Update $\mathcal{Q}:=\mathcal{Q}\cup\{(\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}^{*},\mathcal{F}_{\rm ct}^{*})\}$
21:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
The local alignment of road segments, in Algorithm 7, is established by combining Algorithms 26. Its inputs are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges for the reference linestrings $k$, the blend tolerance $\varepsilon$ and the snap tolerance $\varepsilon_{S}$. In Step 1, we node split the flow linestrings $\mathcal{F}$. Step 2, we node snap the flow linestrings. In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$. In Steps 5–8, we iterate over each reference linestring. In Step 6, we update the reference and candidate flow linestrings by blending the candidate linestrings. In Step 7, we apply ST_OVERLINE_PLANR within each of the $k$-edges of the updated reference linestrings. In Step 8, we compute the weighted mean flow, weighted by the length of the corresponding sub-segments, and assign it as the flow value to all sub-segments. The result from Steps 6–8 is a modified $k$-edge reference linestring with $k$ aggregated flows. In Step 9, we snap the candidate-touching linestrings to the reference linestring to maintain connectivity. In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.
Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align road segment flows using line blending
@@
5:$(\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}):={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
6:$\Edgessubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edges}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
@@
8:     $({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i}^{*}):={\tt ST\_LINEBLEND}({\bm{f}}_{{\rm ref},i},\mathcal{F}_{{\rm cand},i},\varepsilon)$
@@
10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edges}({\bm{f}}_{{\rm ref},i}^{*}))$
@@
We outline the overall workflow ST_OVERLINE to compute a minimal flow map, by combining pre- and post-processing with the iteration of local alignment. For Algorithm 8, the inputs are the map matched routes $\mathcal{M}$, the split node type $S$, the blend tolerance $\varepsilon$, the simplify tolerance $\varepsilon_{D}$, the snap tolerance $\varepsilon_{S}$, the number of connected edges in reference linestrings ${\bm{k}}$, and the maximum number of iterations $j_{\max}$. In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce an initial flow map $\mathcal{F}$. In Step 2, we apply ST_SPLITNODE to this flow map, because applying ST_SPLITNODE on a flow map is more robust than on the map matched routes. In Step 3, we employ the standard ST_SIMPLIFY with simplify tolerance $\varepsilon_{D}$. These simplified linestrings are modified so that all modified segments are at most $\varepsilon_{D}$ distance from the unmodified segments (Ramer, 1972, Douglas and Peucker, 1973). These simplified linestrings usually lead to more overlapping segments, which assist the flow aggregation in Step 4. Steps 5–10 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings. For each $k$ in ${\bm{k}}$, we iterate until the maximum number of iterations $j_{\max}$ is reached or two consecutive flow maps are identical. Within each iteration, we search for the $k$-edge reference linestrings, blend the candidate linestrings, and snap the candidate-touching linestrings, and compute the flow aggregation. In Steps 11–14 is some housekeeping in ST_OVERLINE_PRUNE, where we remove pseudo nodes, and replace the incident edges with a concatenated edge with the weighted mean flow, as well as some loops and leaf edges with low flow values.
Algorithm 8 ST_OVERLINE – Compute locally aligned flow map from map matched routes
1:Input: $\mathcal{M}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}$ aligned road segment flows
3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{M})$;
@@
7:for $k$ in ${\bm{k}}$ do

### v5

@@ 2.4 Line blending to align similar linestrings with local reference
We generalise the reference linestring to be a sequence of $k$ connected edges, each with their own flow. If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings. Due to computational limitations, we retain that candidates be single edges with single flow values. The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
@@
We have described blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges). The next step is to determine the priority for line blending for a set of reference linestrings. We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only. These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them. The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except when $f_{1}=f_{2}$. In this case, then we designate the longer linestring to be the reference.
@@
The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$. The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings. In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e., all paths composed of connected, unique edges) with $k$-edges. In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path. The function ${\rm Edge}(\cdot)$ extracts the component edges from a $k$-edge path. In Step 5, we sort the flow linestrings, in descending order of their flow values and length. In Steps 6–8, we construct a maximal set of non-overlapping $k$-edge paths $\mathcal{K}^{*}$. We initialise $\mathcal{K}^{*}$ to contain the first path. We iterate through $\mathcal{K}$ and, if this path in $\mathcal{K}$ does not overlap the current $\mathcal{K}^{*}$, then we add it to $\mathcal{K}^{*}$. For $k=1$, we bypass Steps 2–8. In Steps 9–18, we step through the ${\bm{k}}_{i}^{*}\in\mathcal{K}^{*}$, starting from the linestring with the highest flow and length. In Steps 11–13, if the current flow linestring ${\bm{k}}_{i}^{*}$ has no incident edges in the candidate linestrings $\mathcal{F}_{\rm cand}$, then we set it to be a reference linestring. In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1). In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$, and $\mathcal{F}_{\rm ct}$. In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets.
@@
6:     $f_{i}$ := weighted mean of $k$ flows, weight := len(Edge(${\bm{k}}_{i}$))
@@
13:     if (${\rm Edge}({\bm{k}}_{i}^{*})\notin\mathcal{F}_{\rm cand}$then
@@
15:         Search $\mathcal{F}_{{\rm cand}}^{*}:=\{{\bm{f}}\in\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\{{\bm{k}}_{\rm ref}\}\cup\mathcal{F}_{\rm cand})\colon{\bm{f}}\subseteq{\tt ST\_BUFFER}({\bm{k}}_{\rm ref},\varepsilon)\}$
16:         if $\mathcal{F}_{\rm cand}^{*}\neq\{\}$ then
17:              Update $\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm ref}\cup\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm cand}^{*}$
18:              Search $\mathcal{F}_{{\rm ct}}:=\{{\bm{f}}\in\mathcal{F}\backslash\{{\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\}\colon{\tt ST\_TOUCHES}(\mathcal{F}_{{\rm cand}}^{*},{\bm{f}})\}$
19:              Update $\mathcal{Q}:=\mathcal{Q}\cup\{(\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}^{*},\mathcal{F}_{\rm ct}^{*})\}$
20:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
The local alignment of flow segments, in Algorithm 7, is established by combining Algorithms 26. The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges $k$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$. In Step 1, we node split the flow linestrings $\mathcal{F}$. In Step 2, we node snap the flow linestrings. In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$. In Steps 5–8, we iterate over each reference linestring. In Step 6, we update the reference and candidate flow linestrings by blending the candidate linestrings. In Step 7, we apply ST_OVERLINE_PLANR within each of the $k$-edges of the updated reference linestrings. In Step 8, we compute the weighted mean flow, weighted by the length of the corresponding sub-segments, and assign it as the flow value to all sub-segments. The result from Steps 6–8 is a modified $k$-edge reference linestring with $k$ aggregated flows. In Step 9, we snap the candidate-touching linestrings to the reference linestring to maintain connectivity. In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.
Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align segment flows using line blending
@@
5:$\{\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}\}:={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
6:$\Edgesubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
@@
8:     $\{{\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i}^{*}\}:={\tt ST\_LINEBLEND}({\bm{f}}_{{\rm ref},i},\mathcal{F}_{{\rm cand},i},\varepsilon)$
@@
10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edge}({\bm{f}}_{{\rm ref},i}^{*}))$
@@
In Figure 5(a) are the flow segments without any line blending, where, for example, the 3 segments with flow 2, 152 and 163, should be blended into a single segment. In Figure 5(b) are the flow segments after line blending. These 3 separate segments are blended into a segment with flow 317 (pink), which is the sum of their individual flows. In contrast, the segments with flows 8, 12 and 16, forming part of the T-junction, are blended into a segment with a weighted mean of 15. Due to the action of ST_OVERLINE_LINEBLEND, the line blended flow segments in Figure 5(b) are an improvement over Figure 5(a).
Refer to caption Refer to caption
Figure 5: (a) Flow segments without line blending. (b) Flow segments with line blending. Label is traffic flow.
To compute a locally aligned flow map, we combine Algorithms 17 into an initial and an iterative phase. The initial phase ST_OVERLINE_INITIAL is described in Algorithm 8. Its inputs are the map matched routes $\mathcal{R}$, the split node type $S$, and the simplify tolerance $\varepsilon_{D}$. In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce a flow map $\mathcal{F}$. In Step 2, we apply ST_SPLITNODE to this flow map. In Step 3, we employ the standard ST_SIMPLIFY with simplify tolerance $\varepsilon_{D}$. These simplified linestrings are modified so that all modified segments are at most $\varepsilon_{D}$ distance from the unmodified segments (Ramer, 1972, Douglas and Peucker, 1973). These simplified linestrings usually lead to more overlapping segments, which assist the flow aggregation in Step 4.
Algorithm 8 ST_OVERLINE_INITIAL – Compute initial locally aligned flow map
1:Input: $\mathcal{R}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance
2:Output: $\mathcal{F}$ initial aligned flow map
3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$
@@
With an initial flow map on hand, we proceed iteratively towards a locally aligned flow map in Algorithm 9. The inputs of ST_OVERLINE are the initial flow map $\mathcal{F}$, the split node type $S$, the blend tolerance $\varepsilon$, the snap tolerance $\varepsilon_{S}$, the number of connected edges ${\bm{k}}$, and the maximum number of iterations $j_{\max}$. Steps 1–6 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings. For each $k$ in ${\bm{k}}$, we iterate until $j_{\max}$ is reached or two consecutive flow maps are identical. Within each iteration, we search for the $k$-edge reference linestrings, blend the candidate linestrings, and snap the candidate-touching linestrings, and compute the flow aggregation. In Steps 7–10 is some housekeeping in ST_OVERLINE_PRUNE, where we remove pseudo nodes, remove leaf edges with low flow values, and replace thin loops with non-looped linestrings.
Algorithm 9 ST_OVERLINE – Compute iterated locally aligned flow map
1:Input: $\mathcal{F}$ initial flow map, $S$ split node, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
2:for $k$ in ${\bm{k}}$ do
3:     $\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
4:     while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
5:         $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
6:         $\mathcal{F}:={\tt ST\_OVERLINE\_LINEBLEND}(\mathcal{F},k,\varepsilon,\varepsilon_{S})$
7:         $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$